Solve the system, using substitution. Write the solution as an ordered pair leaving no spaces, or write no solution or infinitely many solutions; Example: (2,1)
3a - b = 11 and 2a + 3b = 0

Answer by rfer(16322)

-b=-3a+11
b=3a-11
2a+3(3a-11)=0
2a+9a-33=0
11a=33
a=3
-----------------
3(3)-b=11
9-b=11
-b=2
b=-2
------------
3(3)-(-2)=11
(3,-2)